the post office building in manilais 21.8 meters away from the southern bank of the pasig river. from a window of the building 18.5 meters above the ground, a man finds that the angle of depression of the northern bank is 8 degrees and 12 minutes. what is the width of the river at this point?
Answer by lwsshak3(11628)

he post office building in manilais 21.8 meters away from the southern bank of the pasig river. from a window of the building 18.5 meters above the ground, a man finds that the angle of depression of the northern bank is 8 degrees and 12 minutes. what is the width of the river at this point?
***
let x=width of river
8 deg and 12 min=8.2˚
tan(90-8.2)=(21.8+x)/18.5
21.8+x=18.5*tan(81.8˚)=128.38
x=106.58
width of river=106.58 m